OpenStudy (anonymous):

x^2+2x-5 complete the sq

OpenStudy (anonymous):

okay so its (x^2+2/2)^2 for constant apply formula c-1/4*b^2 gives constant = -6 (x^2+1)^2 -6 = ans

OpenStudy (anonymous):

correction (x+1)^2 -6